# Distribution Backtracking Builds A Faster Convergence Trajectory for One-step Diffusion Distillation

### v1

@@ Abstract
Accelerating the sampling speed of diffusion models remains a significant challenge. Recent score distillation methods distill a heavy teacher model into an one-step student generator, which is optimized by calculating the difference between the two score functions on the samples generated by the student model. However, there is a score mismatch issue in the early stage of the distillation process, because existing methods mainly focus on using the endpoint of pre-trained diffusion models as teacher models, overlooking the importance of the convergence trajectory between the student generator and the teacher model. To address this issue, we extend the score distillation process by introducing the entire convergence trajectory of teacher models and propose Distribution Backtracking Distillation (DisBack) for distilling student generators. DisBask is composed of two stages: Degradation Recording and Distribution Backtracking. Degradation Recording is designed to obtain the convergence trajectory of teacher models, which records the degradation path from the trained teacher model to the untrained initial student generator. The degradation path implicitly represents the intermediate distributions of teacher models. Then Distribution Backtracking trains a student generator to backtrack the intermediate distributions for approximating the convergence trajectory of teacher models. Extensive experiments show that DisBack achieves faster and better convergence than the existing distillation method and accomplishes comparable generation performance. Notably, DisBack is easy to implement and can be generalized to existing distillation methods to boost performance. Our code is publicly available on https://github.com/SYZhang0805/DisBack.
Figure 1: The comparison of the distillation process between existing SOTA score distillation method Diff-instruct (Luo et al., 2023b) and our proposed DisBack on (a) CIFAR10, (b) FFHQ64, and (c) ImageNet64 datasets. DisBack achieves faster convergence speed due to the introduction of the entire convergence trajectory between the teacher model and the student generator.
@@ 1 Introduction
Recently, a plethora of generative models have demonstrated remarkable performance across diverse domains such as images (Kou et al., 2023; Yin et al., 2024a), audio (Evans et al., 2024; Xing et al., 2024), and videos (Wang et al., 2024; Chen et al., 2024). However, existing models still grapple with the “trilemma” problem, wherein they struggle to simultaneously achieve high generation quality, fast generation speed, and high sample diversity (Xiao et al., 2021). Generative Adversarial Networks (GANs) (Goodfellow et al., 2014) can rapidly produce high-quality samples but often face mode collapse issues. Variational Autoencoders (VAEs) (Kingma & Welling, 2014) offer stable training but tend to yield lower-quality samples. Recently, Diffusion models (DMs) have emerged as a competitive contender in the generative model landscape (Fan et al., 2023; Zhou et al., 2023; Xu et al., 2024). DMs can generate high-quality, diverse samples but still suffer from slow sampling speeds due to iterative network evaluations.
Figure 2: Several examples of 1024$\times$1024 images generated by our proposed one-step DisBack model distilled from SDXL (Podell et al., ).
@@
To accelerate the sampling speed, the score distillation method tries to distill a large teacher model to a smaller one-step student generator to reduce the sampling cost (Bao et al., 2023; Luo et al., 2023b; Yin et al., 2024b), it optimizes the student model by calculating the difference between two score functions on the samples generated by the student. However, initially as the generated distribution are far from the training distrbution, the generated sample lies outside the training data distribution. Thus, the predicted score of the generated sample from the teacher model does not match the sample’s real score in the training distribution. This mismatch issue is reflected by unreliable network predictions of the teacher model, which prevents the student model from receiving accurate guidance and leads to a decline in final generative performance. We identified that this issue arises because existing score distillation models mainly focus on using the endpoint of the pre-trained diffusion model as the teacher model, overlooking the importance of the convergence trajectory between the student generator and the teacher model. Without the constraint of a convergence trajectory, the mismatch issue causes the student generator to deviate from a reasonable optimization path during training, leading to convergence to suboptimal solutions and a decline in final performance.
@@
To address this problem, we extend the score distillation process by introducing the entire distillation process of the teacher model and propose Distribution Backtracking Distillation (DisBack) for a faster and more efficient distillation. The construction of DisBack is based on the following insights. In practice, the convergence trajectory of most teacher models is inaccessible, particularly for large models like Stable Diffusion (Rombach et al., 2022). Reversely, it is possible to construct a degradation path from the teacher models to the initial student generators, and then the reverse of this path is used as the convergence trajectory of the teacher models. Compared with fitting the teacher model directly, fitting intermediate targets along the convergence trajectory can mitigate the mismatch issue. Thus, the DisBack incorporates degradation recording and distribution backtracking stages. In the degradation recording stage, the teacher model is tuned to fit the distribution of the initial student generator and obtains a distribution degradation path. The path includes a series of in-between diffusion models to represent the intermediate distributions of the teacher model implicitly. In the distribution backtracking stage, the degradation path is reversed and viewed as the convergence trajectory. Then the student generator is trained to backtrack the intermediate distributions along the trajectory to optimize toward the convergence trajectory of the teacher model. In practice, the degradation recording stage typically requires only a few hundred iterations. Therefore, the proposed method incurs trivial additional computational costs. Compared to the existing score distillation method, DisBack exhibits a significantly increased convergence speed (Fig. 1), and it also delivers superior generation performance (Fig. 2).
@@
Our main contributions are summarized as follows. (1) We extend the score distillation process by introducing the entire convergence trajectory of the teacher models and propose Distribution Backtracking Distillation (DisBack), which achieves a faster and more efficient distillation (Sec. 4). (2) Extensive experiments demonstrate that DisBack accelerates the convergence speed of the score distillation process while achieving comparable generation quality compared to existing methods (Sec. 5). (3) The contribution of DisBack is orthogonal to those of other distillation methods. Researchers are encouraged to incorporate our DisBack training strategy into their distillation methods.
@@ 2 Related Works
To improve the efficiency of the diffusion model, existing methods use the knowledge distillation method to distill a large teacher diffusion model to a small and efficient student diffusion model (Yang et al., 2022). The progressive distillation model (Salimans & Ho, 2021) progressively distills the entire sampling process into a new diffusion model with half the number of steps iteratively. Building on this, the classifier-guided distillation model (Sun et al., 2023) introduces a dataset-independent classifier to focus the student model on the crucial features to enhance the distillation process. Guided-distillation model (Meng et al., 2023) proposes a classifier-free guiding framework to avoid the computational cost of additional classifiers and achieve high-quality sampling in only 2-4 steps. Recently, Consistency Model (Song et al., 2023) uses the self-consistency of the ODE generation process to achieve one-step distillation, but this is at the expense of generation quality. To mitigate the surface of the sample quality caused by the acceleration, Consistency Trajectory Models (Kim et al., 2024) combines the adversarial training and denoising score matching loss to further improve the performance. Latent Adversarial Diffusion Distillation (Sauer et al., 2024) leverages generative features from pre-trained latent diffusion models to achieve high-resolution, multi-aspect ratio, few-step image generation.
@@
Diff-Instruct (Luo et al., 2023b) proposes a distillation method from the pre-trained diffusion model to the one-step generator that involves optimizing the generator by the gradient of the difference between two score functions. One score function represents the pre-trained diffusion distribution, while the other represents the generated distribution. Adversarial Score Distillation (Wei et al., 2024) further employs the paradigm of WGAN and retains an optimizable discriminator to improve performance. Additionally, Swiftbrush (Nguyen & Tran, 2024) leverages score distillation to distill a Stable Diffusion v2.1 into a one-step text-to-image generation model and achieve competitive results. DMD (Yin et al., 2024b) suggests the inclusion of a regression loss between noisy images and corresponding outputs to alleviate instability in the distillation process in text-to-image generation tasks. DMD2 (Yin et al., 2024a) introduces a two-time-scale update rule and an additional GAN loss to address the issue of generation quality being limited by the teacher model in DMD, achieving superior performance. Recently, HyperSD (Ren et al., 2024) integrate score distillation with trajectory segmented consistency distillation and the human feedback learning, which achieves SOTA performance from 1 to 8 inference steps.
@@ 3 Preliminary
$\min_{\eta}E_{t\sim\mathcal{U}(0,1),\epsilon\sim\mathcal{N}(0,I)}D_{KL}\left(q_{t}^{G}\left(\bm{x}_{t}\right)\|q_{t}\left(\bm{x}_{t}\right)\right)$ (2)
Here $\bm{x}_{t}$ is the noisy data and $p(\bm{x}_{t}\mid\bm{x}_{0})\sim\mathcal{N}(\bm{x}_{0},\sigma_{t}^{2}I)$. Theorem 1 proves that introducing the additional KL-divergence for $t>0$ does not affect the global optimum of the original optimization problem in Eq.(1).
@@
$\nabla_{\eta}D_{\mathrm{KL}}\left(q^{G}_{t}\left(\bm{x}_{t}\right)\|q_{t}\left(\bm{x}_{t}\right)\right)=E_{t,\epsilon}\left[\left[\nabla_{\bm{x}_{t}}\log q^{G}_{t}\left(\bm{x}_{t}\right)-\nabla_{\bm{x}_{t}}\log q_{t}\left(\bm{x}_{t}\right)\right]\frac{\partial\bm{x}_{t}}{\partial\eta}\right]$ (4)
Here the score of perturbed training data $\nabla_{\bm{x}_{t}}\log q_{t}\left(\bm{x}_{t}\right)$ can be approximated by a pre-trained model $s_{\theta}$. The score of perturbed generated data $\nabla_{\bm{x}_{t}}\log q^{G}_{t}\left(\bm{x}_{t}\right)$ is estimated by another diffusion model $s_{\phi}$, which is optimized by the traditional diffusion training with generated data (Song et al., 2021b):
$\min_{\phi}E_{t,\epsilon}\left\|s_{\phi}\left(\bm{x}_{t},t\right)-\frac{\bm{x}_{0}-\bm{x}_{t}}{\sigma_{t}^{2}}\right\|_{2}^{2}$ (5)
@@
$\nabla_{\eta}D_{\mathrm{KL}}\left(q^{G}_{t}\left(\bm{x}_{t}\right)\|q_{t}\left(\bm{x}_{t}\right)\right)\approx E_{t,\epsilon}\left[s_{\phi}\left(\bm{x}_{t},t\right)-s_{\theta}\left(\bm{x}_{t},t\right)\right]\frac{\partial\bm{x}_{t}}{\partial\eta}$ (6)
@@
In practice, $s_{\phi}$ has three initialization strategies: (1) $s_{\phi}$ is randomly initialized (Franceschi et al., 2023). (2) $s_{\phi}$ is initialized as $s_{\theta}$ or its LoRA (Hu et al., 2021; Wei et al., 2024). (3) $s_{\phi}$ is initialized by fitting the generated samples of student generator (Luo et al., 2023b). Beyond unconditional image generation (Ye & Liu, 2023), this method has also been applied to tasks such as text-to-image and image-to-image generation across various structures (Yin et al., 2024b; Hertz et al., 2023).
@@ 4.1 Insight
In this section, we introduce the Distribution Backtracking Distillation (DisBack). The key insight behind DisBack is the importance of the convergence trajectory. In score distillation, there are two score functions, one representing the pretrained diffusion distribution and the other representing the generated distribution. The student model is optimized using the gradient of the difference between these two score functions. Existing methods (Luo et al., 2023b; Yin et al., 2024b; a) only consider the endpoint of pre-trained diffusion models as teacher models, overlooking the intermediate convergence trajectory between the student generator and the teacher model. The resulting score mismatch issue between the predicted scores of the teacher model and the real scores causes the student model to receive inaccurate guidance and ultimately leads to a decline in final performance. Constraining the convergence trajectory during distillation can mitigate the mismatch issue and help the student generator approximate the convergence trajectory of teacher models to achieve faster convergence. In practice, it is infeasible to obtain the convergence trajectory of most teacher models, especially for large models such as Stable Diffusion. Because the trajectory of distribution changes is bidirectional, it is possible to obtain the degradation path from the teacher model to the initial student generator. The reverse of this degradation path can be regarded as the convergence trajectory of the teacher model. Based on the above insights, we structure the proposed DisBack method in two stages including the degradation recording stage and the distribution backtracking stage (Fig. 3).
@@ 4.2 Degradation Recording
This stage aims to obtain the degradation path from the teacher model to the initial student generator. The degradation path is then reversed and viewed as the convergence trajectory between the student generator and the teacher model. The teacher model here is the pre-trained diffusion model $s_{\theta}$ and the student generator is represented by $G^{0}_{stu}$.
Let $s_{\theta}^{\prime}$ be a diffusion model initialized with the teacher model $s_{\theta}$, and it is trained on generated samples to fit the initial student generator’s distribution $q^{G}_{0}$ with Eq. (7). In this case, the well-trained $s_{\theta}^{\prime}$ is degraded to the untrained distribution $q^{G}_{0}$. By saving the multiple intermediate checkpoints during the training, we can obtain a series of diffusion models $\{s_{\theta_{i}}^{\prime}\mid i=0,\ldots,N\}$, where $s_{\theta_{0}}^{\prime}=s_{\theta}\approx q_{0}$ and $s_{\theta_{N}}^{\prime}\approx q^{G}_{0}$. These diffusion models describe the scores of non-existent distributions on the path recording how the training distribution $q_{0}$ degrades to $q^{G}_{0}$. Algorithm 1 shows the process of obtaining the degradation path. Since distribution degradation is easily achievable, the degradation recording stage only needs trivial additional computational resources (200 iterations in most cases).
$\min_{\theta}E_{t,\epsilon}\left\|s_{\theta}^{\prime}\left(\bm{x}_{t},t\right)-\frac{\bm{x}_{0}-\bm{x}_{t}}{\sigma_{t}^{2}}\right\|_{2}^{2}$ (7)
@@
       $\frac{\partial}{\partial\theta_{i}}E_{t,\epsilon}\left\|s_{\theta}^{\prime}\left(\bm{x}_{t},t\right)-\frac{\bm{x}_{0}-\bm{x}_{t}}{\sigma_{t}^{2}}\right\|_{2}^{2}$
@@
          $E_{t,\epsilon}\left[s_{{\phi}}\left(\bm{x}_{t},t\right)-s_{\theta_{i}}^{\prime}\left(\bm{x}_{t},t\right)\right]\frac{\partial\bm{x}_{t}}{\partial\eta}$
@@
          $\frac{\partial}{\partial\phi}E_{t,\epsilon}\left\|s_{\phi}\left(\bm{x}_{t},t\right)-\frac{\bm{x}_{0}-\bm{x}_{t}}{\sigma_{t}^{2}}\right\|_{2}^{2}$

### v3

# Distribution Backtracking Builds A Faster Convergence Trajectory for Diffusion Distillation

@@ Abstract
Accelerating the sampling speed of diffusion models remains a significant challenge. Recent score distillation methods distill a heavy teacher model into a student generator to achieve one-step generation, which is optimized by calculating the difference between two score functions on the samples generated by the student model. However, there is a score mismatch issue in the early stage of the score distillation process, since existing methods mainly focus on using the endpoint of pre-trained diffusion models as teacher models, overlooking the importance of the convergence trajectory between the student generator and the teacher model. To address this issue, we extend the score distillation process by introducing the entire convergence trajectory of the teacher model and propose Distribution Backtracking Distillation (DisBack). DisBask is composed of two stages: Degradation Recording and Distribution Backtracking. Degradation Recording is designed to obtain the convergence trajectory by recording the degradation path from the pre-trained teacher model to the untrained student generator. The degradation path implicitly represents the intermediate distributions between the teacher and the student, and its reverse can be viewed as the convergence trajectory from the student generator to the teacher model. Then Distribution Backtracking trains the student generator to backtrack the intermediate distributions along the path to approximate the convergence trajectory of the teacher model. Extensive experiments show that DisBack achieves faster and better convergence than the existing distillation method and achieves comparable or better generation performance, with an FID score of 1.38 on the ImageNet 64$\times$64 dataset. DisBack is easy to implement and can be generalized to existing distillation methods to boost performance. Our code is publicly available on https://github.com/SYZhang0805/DisBack.
Figure 1: The comparison of the distillation process between existing SOTA score distillation method Diff-Instruct (Luo et al., 2023c) and proposed DisBack on (a) CIFAR10, (b) FFHQ 64x64, and (c) ImageNet 64x64 datasets. The first 200 epochs refer to the computational overhead of the degradation recording stage of DisBack. DisBack achieves a faster convergence speed due to the constraint of the entire convergence trajectory between the student generator and the teacher model.
@@ 1 Introduction
Recently, generative models have demonstrated remarkable performance across diverse domains such as images (Kou et al., 2023; Yin et al., 2024a), audio (Evans et al., 2024; Xing et al., 2024), and videos (Wang et al., 2024; Chen et al., 2024). However, existing models still grapple with the “trilemma” problem, wherein they struggle to simultaneously achieve high generation quality, fast generation speed, and high sample diversity (Xiao et al., 2021). Generative Adversarial Networks (GANs) (Goodfellow et al., 2014) can rapidly produce high-quality samples but often face mode collapse issues. Variational Autoencoders (VAEs) (Kingma & Welling, 2014) offer stable training but tend to yield lower-quality samples. Recently, Diffusion models (DMs) have emerged as a competitive contender in the generative model landscape (Fan et al., 2023; Zhou et al., 2023; Xu et al., 2024). Diffusion models can generate high-quality, diverse samples but still suffer from slow sampling speeds due to iterative network evaluations.
@@
To accelerate the sampling speed, the score distillation method tries to distill a heavy teacher model to a student generator to reduce the sampling cost and achieve the one-step generation  (Bao et al., 2023; Luo et al., 2023c; Yin et al., 2024b). The score distillation method optimizes the student generator by calculating the difference between two score functions on the samples generated by the student generator. However, as the generated distribution is far from the training distribution at the beginning, the generated sample lies outside the training data distribution. Thus, the predicted score of the generated sample from the teacher model does not match the sample’s real score in the training distribution. This mismatch issue is reflected by unreliable network predictions of the teacher model, which prevents the student model from receiving accurate guidance and leads to a decline in final generative performance. We identified that this issue arises because existing score distillation methods mainly focus on using the endpoint of the pre-trained diffusion model as the teacher model, overlooking the importance of the convergence trajectory between the student generator and the teacher model. Without the constraint of the convergence trajectory, the mismatch issue causes the student generator to deviate from a reasonable optimization path during training, leading to convergence to suboptimal solutions and a decline in final performance.
@@
To address this problem, we extend the score distillation process by introducing the entire convergence trajectory of the teacher model and propose Distribution Backtracking Distillation (DisBack) for a faster and more efficient distillation. The construction of DisBack is based on the following insights. In practice, the convergence trajectory of most teacher models is inaccessible, particularly for large models like Stable Diffusion (Rombach et al., 2022). Because the trajectory of distribution changes is bidirectional, it is possible to construct a degradation path from the teacher model to the initial student generator, and the reverse of this path can be viewed as the convergence trajectory of the teacher model. Compared with fitting the teacher model directly, fitting intermediate targets along the convergence trajectory can mitigate the mismatch issue. Thus, the DisBack incorporates degradation recording and distribution backtracking stages. In the degradation recording stage, the teacher model is tuned to fit the distribution of the initial student generator and obtains a distribution degradation path. The path includes a series of in-between diffusion models to represent the intermediate distributions of the teacher model implicitly. In the distribution backtracking stage, the degradation path is reversed and viewed as the convergence trajectory. Then the student generator is trained to backtrack the intermediate distributions along the path to optimize towards the convergence trajectory of the teacher model. In practice, the degradation recording stage typically requires only a few hundred iterations. Therefore, the proposed method incurs trivial additional computational costs. Compared to the existing score distillation method, DisBack exhibits a significantly increased convergence speed (Fig. 1), and it also delivers superior generation performance (Fig. 2).
@@
Our main contributions are summarized as follows. (1) We extend the score distillation process by introducing the constraint of the entire convergence trajectory of the teacher model and propose Distribution Backtracking Distillation (DisBack), which achieves a faster and more efficient distillation (Sec. 4). (2) Extensive experiments demonstrate that the proposed DisBack accelerates the convergence speed of the score distillation process while achieving comparable or better generation quality compared to existing methods (Sec. 5). (3) The contribution of DisBack is orthogonal to those of other distillation methods. Researchers are encouraged to incorporate our DisBack training strategy into their distillation methods.
Figure 2: Several examples of 1024$\times$1024 images generated by our proposed one-step DisBack model distilled from SDXL (Podell et al., 2024).
@@ 2 Related Works
To improve the efficiency of the diffusion model, existing methods use knowledge distillation to distill a large teacher model to a small and efficient student model (Yang et al., 2022). The progressive distillation (Salimans & Ho, 2021) progressively distills the entire sampling process into a new diffusion model with half the number of steps iteratively. Building on this, the classifier-guided distillation (Sun et al., 2023) introduces a dataset-independent classifier to focus the student on the crucial features to enhance the distillation process. Guided-distillation (Meng et al., 2023) proposes a classifier-free guiding framework to avoid the computational cost of additional classifiers and achieve high-quality sampling in only 2-4 steps. Recently, the Consistency Model (Song et al., 2023) uses the self-consistency of the ODE process to achieve one-step distillation, but this is at the expense of generation quality. To mitigate the surface of the sample quality caused by the acceleration, the Consistency Trajectory Model (Kim et al., 2024a) combines the adversarial training and denoising score matching loss to further improve the performance. Latent Adversarial Diffusion Distillation (Sauer et al., 2024) leverages generative features from pre-trained latent diffusion models to achieve high-resolution, multi-aspect ratio, few-step image generation.
@@
Diff-Instruct (Luo et al., 2023c) distills a pre-trained diffusion model into a one-step generator by optimizing it with the gradient of the difference between two score functions—one for the pre-trained diffusion distribution and the other for the generated distribution. Adversarial Score Distillation (Wei et al., 2024) further employs the paradigm of WGAN and retains an optimizable discriminator to improve performance. Additionally, Swiftbrush (Nguyen & Tran, 2024) leverages score distillation to distill a Stable Diffusion v2.1 into a one-step generator and achieve competitive results. DMD (Yin et al., 2024b) suggests the inclusion of a regression loss between noisy images and corresponding outputs to alleviate instability in the distillation process in text-to-image generation tasks. DMD2 (Yin et al., 2024a) introduces a two-time-scale update rule and an additional GAN loss to address the issue of generation quality being limited by the teacher model in DMD, achieving superior performance. Recently, HyperSD (Ren et al., 2024) integrates score distillation with trajectory segmented consistency distillation and human feedback learning, which achieves SOTA performance from 1 to 8 inference steps.
Mismatch issues in diffusion model.
Existing studies have already discussed the issue of score mismatch in diffusion models Chao et al. (2022); Kim et al. (2024b). DLSM Chao et al. (2022) deeply analyzed the mismatch between the posterior score estimated by diffusion models and the true likelihood score in conditional generation and proposed the Denoising Likelihood Score Matching loss to train classifiers for more accurate conditional score estimation. TIW-DSM Kim et al. (2024b) highlighted that dataset biases cause mismatches between the generator’s and true data distributions, introducing time-varying importance weights and score correction terms to mitigate this issue by assigning different weights to data points at each time step and correcting the scores. This paper focuses on the mismatch arising from the disparity between the initial generator’s distribution and the teacher model’s training distribution, causing discrepancies between predicted and true scores.
@@ 3 Preliminary
$\min_{\eta}\mathbb{E}_{t\sim\mathcal{U}(0,1),\bm{\epsilon}\sim\mathcal{N}(0,I)}D_{KL}\left(q_{t}^{G}\left(\bm{x}_{t}\right)\|q_{t}\left(\bm{x}_{t}\right)\right)$ (2)
Here $\bm{x}_{t}$ is the noisy data and $\bm{x}_{t}\mid\bm{x}_{0}\sim\mathcal{N}(\bm{x}_{0},\sigma_{t}^{2}I)$. Theorem 1 proves that introducing the additional KL-divergence for $t>0$ does not affect the global optimum of Eq.(1)
@@
$\nabla_{\eta}D_{\mathrm{KL}}\left(q^{G}_{t}\left(\bm{x}_{t}\right)\|q_{t}\left(\bm{x}_{t}\right)\right)=\mathbb{E}_{t,\bm{\epsilon}}\left[\left[\nabla_{\bm{x}_{t}}\log q^{G}_{t}\left(\bm{x}_{t}\right)-\nabla_{\bm{x}_{t}}\log q_{t}\left(\bm{x}_{t}\right)\right]\frac{\partial\bm{x}_{t}}{\partial\eta}\right]$ (4)
Here the score of perturbed training data $\nabla_{\bm{x}_{t}}\log q_{t}\left(\bm{x}_{t}\right)$ can be approximated by a pre-trained diffusion model $s_{\theta}$. The score of perturbed generated data $\nabla_{\bm{x}_{t}}\log q^{G}_{t}\left(\bm{x}_{t}\right)$ is estimated by another diffusion model $s_{\phi}$, which is optimized by score matching with generated data (Song et al., 2021b):
$\min_{\phi}\mathbb{E}_{t,\bm{\epsilon}}\left\|s_{\phi}\left(\bm{x}_{t},t\right)-\frac{\bm{x}_{0}-\bm{x}_{t}}{\sigma_{t}^{2}}\right\|_{2}^{2}$ (5)
@@
$\nabla_{\eta}D_{\mathrm{KL}}\left(q^{G}_{t}\left(\bm{x}_{t}\right)\|q_{t}\left(\bm{x}_{t}\right)\right)\approx\mathbb{E}_{t,\bm{\epsilon}}\left[\left[s_{\phi}\left(\bm{x}_{t},t\right)-s_{\theta}\left(\bm{x}_{t},t\right)\right]\frac{\partial\bm{x}_{t}}{\partial\eta}\right]$ (6)
@@
In practice, $s_{\phi}$ has three initialization strategies: (1) $s_{\phi}$ is randomly initialized (Franceschi et al., 2023). (2) $s_{\phi}$ is initialized as $s_{\theta}$ or its LoRA (Hu et al., 2021; Wei et al., 2024). (3) $s_{\phi}$ is initialized by fitting the generated samples of student generator (Luo et al., 2023c). Beyond unconditional image generation (Ye & Liu, 2023), this method has also been applied to tasks such as text-to-image and image-to-image generation across various structures (Yin et al., 2024b; Hertz et al., 2023).
@@ 4.1 Insight
In this section, we introduce the Distribution Backtracking Distillation (DisBack). The key insight behind DisBack is the importance of the convergence trajectory. As mentioned in Sec.3, there are two score functions in score distillation, one representing the pre-trained diffusion distribution and the other representing the generated distribution. The student model is optimized using the gradient of the difference between these two score functions. Existing methods (Luo et al., 2023c; Yin et al., 2024b; a) directly use the endpoint of the pre-trained diffusion model as the teacher model, overlooking the intermediate convergence trajectory between the student generator and the teacher model. The resulting score mismatch issue between the predicted scores of the generated sample from the teacher model and the real scores causes the student model to receive inaccurate guidance. It ultimately leads to a decline in final performance. Constraining the convergence trajectory between the student generator and the teacher model during the distillation process can mitigate the mismatch issue and help the student generator approximate the convergence trajectory of teacher models to achieve faster convergence. In practice, it is infeasible to obtain the convergence trajectory of most teacher models, especially for large models such as Stable Diffusion (Rombach et al., 2022). Reversely, it is possible to obtain the degradation path from the teacher model to the initial student generator. The reverse of this degradation path can be viewed as the convergence trajectory of the teacher model. Based on the above insights, we structure the proposed DisBack in two stages including the degradation recording stage and the distribution backtracking stage (Fig. 3).
@@ 4.2 Degradation Recording
This stage aims to obtain the degradation path from the teacher to the initial student generator. The degradation path is then reversed and viewed as the convergence trajectory of the teacher model. The teacher here is the pre-trained diffusion model $s_{\theta}$ and the student is represented by $G^{0}_{stu}$.
Let $s_{\theta}^{\prime}$ be a diffusion model initialized with the teacher model $s_{\theta}$, and it is trained on generated samples to fit the initial student generator’s distribution $q^{G}_{0}$ with Eq. (7). By saving the multiple intermediate checkpoints during the training, we can obtain a series of diffusion models $\{s_{\theta_{i}}^{\prime}\mid i=0,\ldots,N\}$, where $s_{\theta_{0}}^{\prime}=s_{\theta}\approx q_{0}$ and $s_{\theta_{N}}^{\prime}\approx q^{G}_{0}$. These diffusion models describe the scores of non-existent distributions on the path, recording how the training distribution $q_{0}$ degrades to the initial generated distribution $q^{G}_{0}$. Algorithm 1 shows the process of obtaining the degradation path. Since distribution degradation is easily achievable, the degradation recording stage only needs trivial additional computational resources (200 iterations in most cases).
$\min_{\theta}\mathbb{E}_{t,\bm{\epsilon}}\left\|s_{\theta}^{\prime}\left(\bm{x}_{t},t\right)-\frac{\bm{x}_{0}-\bm{x}_{t}}{\sigma_{t}^{2}}\right\|_{2}^{2}$ (7)
@@
       $\frac{\partial}{\partial\theta_{i}}\mathbb{E}_{t,\bm{\epsilon}}\left\|s_{\theta}^{\prime}\left(\bm{x}_{t},t\right)-\frac{\bm{x}_{0}-\bm{x}_{t}}{\sigma_{t}^{2}}\right\|_{2}^{2}$
@@
          $\mathbb{E}_{t,\bm{\epsilon}}\left[s_{{\phi}}\left(\bm{x}_{t},t\right)-s_{\theta_{i}}^{\prime}\left(\bm{x}_{t},t\right)\right]\frac{\partial\bm{x}_{t}}{\partial\eta}$
@@
          $\frac{\partial}{\partial\phi}\mathbb{E}_{t,\bm{\epsilon}}\left\|s_{\phi}\left(\bm{x}_{t},t\right)-\frac{\bm{x}_{0}-\bm{x}_{t}}{\sigma_{t}^{2}}\right\|_{2}^{2}$